How much interest will P8,500 earn after a year at 7% interest per annum? How much is the total amount due?
Answer by stanbon(75887)

How much interest will P8,500 earn after a year at 7% interest per annum? How much is the total amount due?
----
Interest = 0.07(8500)= $595
Total amount = 8500+595=$9,095
Cheers,
Stan H.